How old am if 500 reduced by 3 times my age is 257?
Answer by ankor@dixie-net.com(22740)

How old am if 500 reduced by 3 times my age is 257?
let a = your age
:
500-3a = 257
500 - 257 = 3a
243 = 3a
a = 243/3
a = 81 is your age